OpenStudy (anonymous):

Which polynomial identity will prove that 117 = 125 - 8?

OpenStudy (anonymous):

subtraction duh......

OpenStudy (anonymous):

Difference of Cubes Difference of Squares Square of Binomial Sum of Cubes

OpenStudy (anonymous):

difference of cubes?

OpenStudy (anonymous):

can i ask a different question ?

OpenStudy (anonymous):

Given a polynomial f(x), if (x + 2) is a factor, what else must be true?

OpenStudy (solomonzelman):

\(\normalsize\color{black}{117 = 125 - 8}\) \(\normalsize\color{black}{117 = 5^3 - 2^3}\) \(\normalsize\color{black}{117 = (5-2)(5^2+(3)(5)+2^2)}\) \(\normalsize\color{black}{117 = (5-2)(25+15+4)}\) \(\normalsize\color{black}{117 = (3)(44)}\) \(\normalsize\color{black}{117 = 132}\) LOL, where is the mistake ?

OpenStudy (solomonzelman):

your options for the next question?

OpenStudy (anonymous):

what in the heck?

OpenStudy (solomonzelman):

IT is certainly true that there is another factor and it involves a variable

OpenStudy (anonymous):

how did you do it sooo fast?

OpenStudy (anonymous):

f(0) = 2 f(0) = -2 f(2) = 0 f(-2) = 0

OpenStudy (solomonzelman):

f(x) = (x+2) × (other factor) f(-2) = (-2+2) × (other factor) f(-2) = 0 × (other factor) f(-2) = 0 what do you think, now ?

OpenStudy (solomonzelman):

I'm not

OpenStudy (solomonzelman):

Shouldn't it make sense, `I'mn't ` ?

OpenStudy (anonymous):

@SolomonZelman and @am!rah thanks

OpenStudy (solomonzelman):

yw !